### v1

@@ 4.3 Model uncertainty
For uncertainty estimation we employ Monte Carlo (MC) Dropout [gal2016dropoutbayesianapproximationrepresenting] as a computationally efficient method to perform probabilistic inference and estimate the predictive uncertainty of our deep learning model. In this context, it is useful to distinguish between two primary types of uncertainty [kendall2017uncertaintiesneedbayesiandeep]: epistemic uncertainty, which captures our ignorance about the model parameters and can be reduced with more data, and aleatoric uncertainty, which captures the inherent noise or randomness in the data itself. The key idea behind MC Dropout is to retain dropout during inference, not just training. Given an input $\mathbf{x}_{i}$, we perform $T$ stochastic forward passes through the trained network $f(\cdot;\theta)$, applying different dropout masks in each pass. This yields a set of $T$ output samples,
@@
The square root of this variance, $\hat{\sigma}(\hat{y}_{ij})=\sqrt{\widehat{\text{Var}}(\hat{y}_{ij})}$ is the standard deviation for each parameter component, which serves as our primary measure of model uncertainty. The sample variance calculated via MC Dropout serves as an approximation of the total predictive uncertainty. It implicitly combines both the epistemic uncertainty, arising from the stochasticity of the dropout masks, which approximates uncertainty over the model’s weights, and the aleatoric uncertainty that is propagated through the deterministic parts of the network from the input data [kendall2017uncertaintiesneedbayesiandeep]. For a computation of the Bayesian posterior which captures all kinds of uncertainty, an approach such as simulation based inference would be appropriate.
@@ 5 Results
Here, we present our results for the parameter inference of noisy waveforms and overall performance of our model. To understand the computational requirements and training dynamics, we trained our model on a small ($\sim 20,000$), intermediate ($\sim 50,000$), and large ($\sim 80,000$) dataset. The train/validation/testing sets were created using our strategy described in the Section 4.1, whereas the small and intermediate ones taken as subsets of the large training set. For all experiments the validation and test sets remained fixed to ensure that the performance improvements observed are attributable to the increase in training data, not variations in the evaluation sets. We notice that all training and evaluation were performed on a single NVIDIA A100 GPU.
@@ 5.1 Effect of pretraining
The results associated to the effect of pretraining at fixed detector noise (ET) are summarised in Tables  1 - 3. Tables 1 and 2 summarise the effect of pretraining on the convergence speed and performance metric respectively, comparing training from scratch and training/fine-tuning on top of a pretrained model. The relative performance gains are shown in Table 3. For each combination of dataset size and signal-to-noise ratio (SNR), the “Target Loss” corresponds to the best validation loss achieved by a model trained from scratch (no pretraining). The results clearly show that pretraining substantially accelerates convergence across all SNRs and dataset sizes, with speedup factors ranging from approximately $1.4\times$ to an impressive speedup of $6.6\times$. The most pronounced benefits of pretraining are clearly observed for small and intermediate datasets, where convergence can be more than $6\times$ faster, such as for SNR = 10 in the small and intermediate dataset scenarios. Remarkably, the speedup advantage is generally greater at lower SNRs, indicating that pretraining is especially beneficial when the signal is weaker and the learning task is more challenging. For large datasets, the speedup effect is less pronounced, particularly at SNR = 10, though the reduction in training time remains sizable. At higher SNR with a large dataset, the speedup is still notable, suggesting that while having a large amount of data partially compensates for the absence of pretraining, the use of pretrained models continues to offer substantial gains. Overall, pretraining in GraviBERT is highly effective for reducing training times and enables more rapid model development, which is particularly valuable in computationally intensive experiments.
@@
Pretraining enhances also the performance metrics of the model. The absolute performance metrics for the pretrained and fine-tuned models are summarized in Table 2, while the fractional gains compared to training from scratch in Table 3. The metrics are computed as the average over 100 stochastic forward passes per sample using MC Dropouts based on the test set. Across all evaluated cases, the pretrained/fine-tuned model consistently outperforms the model trained from scratch, achieving lower MAE for all parameters and, in some cases, higher $R^{2}$ values although the latter is typically mildly affected. At fixed dataset size, pretraining significantly improves MAE with increasing SNR, with the biggest effect in the large dataset where the percentage MAE improvement is about $3\times$ larger. For the most challenging case, specifically the case of SNR = 10 and the small dataset, the benefits of pretraining are pronounced, with a MAE reduction ranging between $6\%-14\%$ across parameters. The best reduction is found for the case of the large dataset and SNR = 30, ranging between $11\%-30\%$. This can be attributed to the sufficiently large amount of data in the large dataset combined with the high SNR which allow the model to adopt and train better. One notices that the intermediate dataset shows suboptimal gains compared to both smaller and larger datasets. This appears to be due to the dataset being large enough for the model to start overfitting to the training data, but not large enough to fully benefit from the pretrained representations or to learn robust patterns from scratch. Another possibility is that the pretrained features may compete with task-specific learning in a suboptimal way leading to this behavior. Furthermore, this could also be (partially) attributed to the use of a single set of hyperparameters across all experiments. The optimal hyperparameters for fine-tuning a pretrained model, which often benefit from a smaller learning rate, may differ from those for training a model from scratch. A more exhaustive study would involve performing a separate hyperparameter search for each condition. However, our results using a single, tuned configuration consistently demonstrate a network positive benefit from pretraining across all dataset sizes. We notice that for all the above cases $R^{2}\geq 0.90$ as shown in Table 2.
@@
Understanding how model performance scales with dataset size informs future data requirements and expected performance gains  [kaplan2020scalinglawsneurallanguage]. The test loss (measured on models trained from scratch) follows a power-law trend as a function of training dataset size $D$,
@@ 5.2 Performance on parameter inference
Figure 8 shows the precision of the model’s predictions with respect to the target values across our grid. For the component masses, we evaluate the mean relative precision across our predictions, defined as $100\times(\text{prediction}-\text{target})/\text{target}$. For the effective spin, we report the mean of the absolute difference $(\text{prediction}-\text{target})$, in order to avoid amplification due to division by values near zero. The mean relative precision across the sample for $m_{1}$ is $6\%$ at SNR = 10 and $3.3\%$ at SNR = 30. The similar values for $m_{2}$ are $3.4\%$ and $1.1\%$, while for distance $8.3\%$ and $4.1\%$ respectively. For $\chi_{\rm{eff}}$, the absolute mean error value is $3.8\times 10^{-3}$ at SNR = 10 and $6\times 10^{-5}$ at SNR = 30. To enable an approximate comparison with real observations, we focus on model predictions at SNR = 10, and select the event GW190828_065509 from the third Gravitational Wave Transient Catalog (GWTC-3) [Abbott2023gwtc3] of the LIGO-Virgo (LV) Collaboration, which has an SNR compatible with $10$  [^3] The reported central values for masses and effective spin are $m_{1}=23.7M_{\odot},m_{2}=10.4M_{\odot},\chi_{\rm{eff}}=0.05$.. To compare our model’s mean relative errors with the $90\%$ credible intervals reported by the LV Collaboration, we convert them to equivalent $90\%$ errors under the assumption of Gaussianity, which amounts to multiplying them by $\simeq 1.645$. For $\{m_{1},m_{2},\chi_{\rm{eff}},d\}$ we find for the average $90\%$ error across our sample $\{9.9\%,5.7\%,0.006,13.6\%\}$, while the corresponding errors reported by the LV Collaboration for the event GW190828_065509 are $\{28.5\%,28.9\%,0.165,44\%\}$. Notice that we have symmetrised the observation errors to facilitate comparison, i.e $\sigma_{\rm 90\%}=(\sigma_{\rm 90\%upper}+\sigma_{\rm 90\%lower})/2$. These numbers should not be interpreted as a direct posterior-to-posterior comparison, and serves only as a back-of-the-envelope estimate. A Bayesian analysis would be needed for a proper comparison. We also emphasise that our distance refers to an effective luminosity distance rescaled to fixed SNR, not an actual astrophysical one.
@@ 5.3 Domain adaptation
Domain adaptation refers to transferring knowledge across domains and is central to foundation-style models, which learn general rather than domain-specific representations. Effective adaptation indicates that the model has captured fundamental structures in the data, with the pretraining process providing robust, transferable features which subsequently require only light fine-tuning. In our case, this invariance is assessed only within the shared ET–LV frequency band ($\geq 20$ Hz). We focus on transferring from ET to LV noise, evaluating whether the model can adopt detector-agnostic GW features rather than memorising detector-specific noise.
@@
We first evaluate zero-shot inference by applying the model trained purely on ET noise directly to the LV one, without any fine-tuning or weight updates. This results in a loss of predictive power, indicating that ET-trained features do not transfer to the LV domain in a zero-shot setting. This is naturally explained by the model having learned a detector-conditioned latent space rather than an invariant one, and therefore, requiring domain adaptation through fine-tuning. Implementing transfer learning through fine tuning, the performance gain is substantial compared to training from scratch across all inferred parameters, ranging between $9\%-47\%$ improvement for MAE (see Tables 4 and 5). The spin parameter shows the biggest improvements ($21-47\%$), suggesting that the pretrained model learned particularly robust representations of spin-related features that transfer effectively between detectors. The mass parameters and distance also show consistent improvements of $13-27\%$, indicating that the model successfully captures detector-agnostic physical relationships. The convergence speed results show that the pretrained model achieves the same performance targets as models trained from scratch in dramatically less time, up to $15\times$ faster. This speedup is particularly pronounced at lower SNR = 10, where the pretrained model reaches target performance in just 2-3 minutes compared to 30-50 minutes for scratch training. Interestingly, the speedup is somewhat reduced SNR = 30, possibly because cleaner signals allow scratch-trained models to converge more efficiently. These results collectively demonstrate that pretraining on one detector configuration creates transferable representations that not only improve final performance, but also significantly reduce the computational cost of adapting to new detector environments. We remind that while fine-tuning on LV noise shows successful adaptation, both detectors still share broadly similar noise structure above 20 Hz. The more demanding task of adapting LV-trained models to ET’s full 5 Hz range remains open for future work.
@@
To evaluate the generalisation of GraviBERT against waveform modelling, we performed a cross-approximant validation. We first reconstructed the held-out test samples from the large dataset (originally generated with SEOBNRv4) using the IMRPhenomD approximant, adjusting the signals to match the temporal grid of the original SEOBNRv4 training data. Inference was then performed on these new waveforms using the frozen weights of our fully trained model (pretrained and trained on ET noise and SEOBNRv4 approximant) without any further fine-tuning, that is, at zero-shot inference. The results are shown in Table 6. While the baseline model achieves $R^{2}>0.93$ for all parameters at SNR = 10, the cross-approximant inference appears to introduce a systematic bias. This effect is particularly evident for the effective spin parameter, where $R^{2}=0.65$ at SNR = 10. Spin is intimately connected with the subtle phase evolution during inspiral, which is modelled differently in SEOBNRv4 and IMRPhenomD. Nevertheless, the model retains predictive power for the dominant parameters $(m_{1},m_{2},d)$ at zero-shot inference, with $R^{2}$ remaining consistently between [0.80,0.97]. This suggests that the model has extracted representations of the signal envelope and frequency chirp, which are approximant-agnostic, while the specific phase alignment remains sensitive to the training domain.
@@ 5.4 Benchmarking
George & Huerta (2018) [george2018deep2_inference] demonstrated CNN-based parameter estimation achieving mean relative errors below 10% for mass parameters on real LV data. Shen et al. (2022) [shen2022deep] achieved 2 ms inference time using Bayesian neural networks trained on over 10 million waveforms for 4-parameter estimation, with relative errors of 7–12% for masses and 5–13% for remnant properties at SNR $\geq$ 15.
@@ 6 Conclusions and Outlook
We have introduced GraviBERT, a novel and versatile model for parameter inference of GW time-series, through the fusion of an Inception-inspired multi-scale convolutional frontend with a transformer encoder. GraviBERT infers source parameters, such as masses and spins, directly from noisy time-domain waveforms down to few-percent level, without the need for explicit denoising steps. Our two-stage training strategy, leveraging BERT-style pretraining followed by targeted fine-tuning, yields substantial improvements in both training efficiency and inference accuracy, and enables adaptation to new detector configurations and waveform approximants. When evaluated on in-domain data (ET), pretraining reduces the mean absolute error by up to $30\%$ and accelerates training convergence by up to $6\times$. Furthermore, the mean relative precision of the inferred masses and distances reaches the few-percent level, while the mean absolute error in the effective spin is as low as $\sim 10^{-3}$ for the challenging case of SNR = 10  [^4] We notice that these results were derived without a thorough hyperparameter study for our model.. We remind that our distance refers to an effective luminosity distance rescaled to fixed SNR, not an actual astrophysical one.
The advantages are even more pronounced in domain adaptation from ET to LV data. By fine-tuning on a small target dataset on LV data, the pretrained model on ET noise converges up to $15\times$ faster and achieves an error reduction of up to $40\%$ compared to a model trained from scratch on LV. This enhanced predictive precision is particularly notable in the low SNR regime, while the flexibility of the model’s architecture paves the way for broad applicability across multi-messenger astronomy. We note that while our estimation of inherent model uncertainty relied on the probabilistic approach of MC Dropouts, our parameter inference itself was not Bayesian. A Bayesian treatment could be pursued by replacing the regression head with normalising flows.
@@
The significance of developing foundation-style models lies in their ability to generalise across a wide range of tasks and datasets, serving as adaptable machines for diverse downstream applications. Foundation models have transformed fields such as natural language processing enabling rapid fine-tuning and transfer learning. For GW astrophysics, they promise a similar leap: accelerate scientific discovery by enabling fast, accurate, and robust inference as new detectors, noise environments, and physical scenarios emerge. In this regard, GraviBERT represents an important step forward towards establishing a flexible, pretrained baseline framework for multi-messenger astronomy, which can be subsequently fine-tuned efficiently.
Exciting avenues for future work include applying GraviBERT to test subtle deviations from GR in gravitational waveforms, as well as probing new physics—such as dark energy—through its impact on the waveform’s amplitude and phase evolution [Berti2015, Saltas:2018fwy, Abac:2025saz, LISA:2022kgy, Barausse:2020rsu]. A critical next step is to extend our analysis to fully Bayesian parameter estimation, and to enable inference of the complete set of source parameters for binary mergers. Additionally, our modelling can be expanded to cover neutron star mergers, extreme mass ratio inspirals, alternative waveform approximants, and predicting the time to merger from inspiral waveforms. A promising direction is also the extension of GraviBERT to multi-messenger inference by incorporating electromagnetic and neutrino signals. On the architectural front, enhancements such as frequency-domain inputs and comprehensive hyperparameter optimisation could significantly improve performance. Together, these developments will help establish GraviBERT as a robust multi-modal framework for multi-messenger astrophysics.
Acknowledgements
We thank François Charton, Nikos Karnesis, Fillipos Kokkinos, Nick Stergioulas and David Trestini for discussions and feedback. This work was funded by the Czech Grant Agency (GAČR) under the grant number 21-16583M. The bulk of our computations were carried out at the MetaCentrum cluster in Czechia.
@@ Acknowledgements
Table 2: Per-parameter metrics across different dataset sizes and SNR values for the model trained from scratch (not pretrained) and the pretrained and fine-tuned model, evaluated on the held-out test set with size of $10,000$ samples. The same detector configuration is kept throughout (ET). The reported metrics are measured on the mean values obtained from an ensemble of $100$ stochastic forward passes using MC Dropout for each test sample. We remind that $d$ corresponds to an effective luminosity distance rescaled to fixed SNR.
@@
Table 3: Relative MAE improvement from pretraining followed by fine-tuning, compared to training from scratch, according to the results of Table 2. Positive values indicate the percentage reduction in error achieved through pretraining.
@@
Table 4: Domain adaptation performance of the pretrained model evaluated with MAE. The comparison is between the model pretrained on ET noise and fine-tuned on LV noise waveforms, and a model trained on LV noise waveforms from scratch. The clean waveforms used cross both datasets were generated using the same procedure, with only the detector noise differing between domains.
@@
Table 5: Same as Table 4, but comparing the convergence speed between the model pretrained on ET noise and fine-tuned on LV noise waveforms, and a model trained on LV noise waveforms from scratch. As before, the “Target Loss" is the best validation loss achieved by the corresponding model trained from scratch and the speedup factor quantifies how many times faster the pretrained model reaches this same performance target.
@@
Table 6: Zero-shot performance of the model on a new waveform approximant at fixed detector noise (ET), quantified using MAE and $R^{2}$. The model was pretrained and subsequently trained on SEOBNRv4. It can be seen that zero-shot inference (without fine-tuning) already yields encouraging performance in the cross-approximant setting (see also Section 5.3).
@@
Table 7: Domain adaptation performance of the pretrained model on a new waveform approximant at fixed detector noise (ET), evaluated with MAE. The comparison is between the model pretrained on SEOBNRv4 waveforms and fine-tuned on IMRPhenomD waveforms, and a model trained on IMRPhenomD waveforms from scratch.
@@
Table 10: Hyperparameter configuration for the pretraining and training stages of the model.

### v2

@@ 4.3 Model uncertainty
For uncertainty estimation we employ Monte Carlo (MC) Dropout [26] applied on the positional encodings, transformer encoder and the regression head. We use it as a computationally efficient method to perform probabilistic inference and estimate the predictive uncertainty of our deep learning model. In this context, it is useful to distinguish between two primary types of uncertainty [33]: epistemic uncertainty, which captures our ignorance about the model parameters and can be reduced with more data, and aleatoric uncertainty, which captures the inherent noise or randomness in the data itself. The key idea behind MC Dropout is to retain dropout during inference, not just training. Given an input $\mathbf{x}_{i}$, we perform $T$ stochastic forward passes through the trained network $f(\cdot;\theta)$, applying different dropout masks in each pass. This yields a set of $T$ output samples,
@@
The square root of this variance, $\hat{\sigma}(\hat{y}_{ij})=\sqrt{\widehat{\text{Var}}(\hat{y}_{ij})}$ is the standard deviation for each parameter component, which serves as our primary measure of model uncertainty. The sample variance calculated via MC Dropout serves as an approximation of the total predictive uncertainty. It implicitly combines both the epistemic uncertainty, arising from the stochasticity of the dropout masks, which approximates uncertainty over the model’s weights, and the aleatoric uncertainty that is propagated through the deterministic parts of the network from the input data [33]. For a computation of the Bayesian posterior which captures all kinds of uncertainty, an approach such as simulation based inference would be appropriate.
@@ 5 Results
Here, we present our results for the parameter inference of noisy waveforms and overall performance of our model. To understand the computational requirements and training dynamics, we trained our model on a small ($\sim 20,000$), intermediate ($\sim 50,000$), and large ($\sim 80,000$) dataset. The train/validation/testing sets were created using our strategy described in the Section 4.1, whereas the small and intermediate ones taken as subsets of the large training set. For all experiments the validation and test sets remained fixed to ensure that the performance improvements observed are attributable to the increase in training data, not variations in the evaluation sets. We note that all training and evaluation were performed on a single NVIDIA A100 GPU.
@@ 5.1 Effect of pretraining
The results associated to the effect of pretraining at fixed detector noise (ET) are summarised in Tables  1 - 3. Tables 1 and 2 summarise the effect of pretraining on the convergence speed and performance metric respectively, comparing training from scratch and training/fine-tuning on top of a pretrained model. The relative performance gains are shown in Table 3. For each combination of dataset size and signal-to-noise ratio (SNR), the “Target Loss” corresponds to the best validation loss achieved by a model trained from scratch (no pretraining). The results clearly show that pretraining substantially accelerates convergence across all SNRs and dataset sizes, with speedup factors ranging from approximately $1.4\times$ to $6.6\times$. The most pronounced benefits of pretraining are clearly observed for small and intermediate datasets, where convergence can be more than $6\times$ faster, such as for SNR = 10 in the small and intermediate dataset scenarios. Remarkably, the speedup advantage is generally greater at lower SNRs, indicating that pretraining is especially beneficial when the signal is weaker and the learning task is more challenging. For large datasets, the speedup effect is less pronounced, particularly at SNR = 10, though the reduction in training time remains sizable. At higher SNR with a large dataset, the speedup is still notable, suggesting that while having a large amount of data partially compensates for the absence of pretraining, the use of pretrained models continues to offer substantial gains. Overall, pretraining in GraviBERT consistently reduces training times and enables more rapid model development, which is particularly valuable in computationally intensive experiments. The convergence behaviour is illustrated in Figure 1(b). The pretrained model reaches the validation target in shorter time while continuing to improve, finally converging to a lower final loss in all configurations.
@@
Pretraining enhances also the performance metrics of the model. The absolute performance metrics for the pretrained and fine-tuned models are summarised in Table 2, while the fractional gains compared to training from scratch in Table 3. The metrics are computed as the average over 100 stochastic forward passes per sample using MC Dropouts based on the test set. Across all evaluated cases, the pretrained/fine-tuned model consistently outperforms the model trained from scratch, achieving lower MAE for all parameters and, in some cases, higher $R^{2}$ values although the latter is typically mildly affected. At fixed dataset size, pretraining significantly improves MAE with increasing SNR, with the biggest effect in the large dataset where the percentage MAE improvement is about $3\times$ larger. For the most challenging case, specifically the case of SNR = 10 and the small dataset, the benefits of pretraining are pronounced, with a MAE reduction ranging between $6\%-14\%$ across parameters. The best reduction is found for the case of the large dataset and SNR = 30, ranging between $11\%-31\%$. This can be attributed to the sufficiently large amount of data in the large dataset combined with the high SNR which allow the model to adopt and train better. One notices that the intermediate dataset shows suboptimal gains compared to both smaller and larger datasets. This appears to be due to the dataset being large enough for the model to start overfitting to the training data, but not large enough to fully benefit from the pretrained representations or to learn robust patterns from scratch. Another possibility is that the pretrained features may compete with task-specific learning in a suboptimal way leading to this behavior. Furthermore, this could also be (partially) attributed to the use of a single set of hyperparameters across all experiments. The optimal hyperparameters for fine-tuning a pretrained model, which often benefit from a smaller learning rate, may differ from those for training a model from scratch. A more exhaustive study would involve performing a separate hyperparameter search for each condition. However, our results using a single, tuned configuration consistently demonstrate a network positive benefit from pretraining across all dataset sizes. We notice that for all the above cases $R^{2}\geq 0.90$ as shown in Table 2.
@@
Understanding how model performance scales with dataset size informs future data requirements and expected performance gains  [32]. The test loss (measured on models trained from scratch) follows a power-law trend as a function of training dataset size $D$,
@@ 5.2 Performance on parameter inference
Figure 8 shows the precision of the model’s predictions with respect to the target values across our grid. For the component masses, we evaluate the mean relative precision across our predictions, defined as $100\times(\text{prediction}-\text{target})/\text{target}$. For the effective spin, we report the mean of the absolute difference $(\text{prediction}-\text{target})$, in order to avoid amplification due to division by values near zero. The mean relative precision across the sample for $m_{1}$ is $6\%$ at SNR = 10 and $3.3\%$ at SNR = 30. The similar values for $m_{2}$ are $3.4\%$ and $1.1\%$, while for distance $8.3\%$ and $4.1\%$ respectively. For $\chi_{\rm{eff}}$, the absolute mean error value is $3.8\times 10^{-3}$ at SNR = 10 and $6\times 10^{-5}$ at SNR = 30. We note that in both cases the predictions are averaged within each grid point group, effectively marginalising over the remaining parameters. To enable an approximate comparison with real observations, we focus on model predictions at SNR = 10, and select the event GW190828_065509 from the third Gravitational Wave Transient Catalog (GWTC-3) [3] of the LIGO-Virgo-Kagra (LVK) Collaboration, which has an SNR compatible with $10$  [^4] The reported central values for masses and effective spin are $m_{1}=23.7M_{\odot},m_{2}=10.4M_{\odot},\chi_{\rm{eff}}=0.05$.. To compare our model’s mean relative errors with the $90\%$ credible intervals reported by the LVK Collaboration, we convert them to equivalent $90\%$ errors under the assumption of Gaussianity, which amounts to multiplying them by $\simeq 1.645$. For $\{m_{1},m_{2},\chi_{\rm{eff}},d\}$ we find for the average $90\%$ error across our sample $\{9.9\%,5.7\%,0.006,13.6\%\}$, while the corresponding errors reported by the LVK Collaboration for the event GW190828_065509 are $\{28.5\%,28.9\%,0.165,44\%\}$. Notice that we have symmetrised the observation errors to facilitate comparison, i.e $\sigma_{\rm 90\%}=(\sigma_{\rm 90\%upper}+\sigma_{\rm 90\%lower})/2$. These numbers should not be interpreted as a direct posterior-to-posterior comparison, and serves only as a back-of-the-envelope estimate. A Bayesian analysis would be needed for a proper comparison. We also emphasise that our distance refers to an effective luminosity distance rescaled to fixed SNR, not an actual astrophysical one.
@@ 5.3 Domain adaptation
Domain adaptation refers to transferring knowledge across domains and is central to foundation-style models, which learn general rather than domain-specific representations. Effective adaptation indicates that the model has captured fundamental structures in the data, with the pretraining process providing robust, transferable features which subsequently require only light fine-tuning. In our case, this invariance is assessed only within the shared ET–LVK frequency band ($\geq 20$ Hz). We focus on transferring from ET to LVK noise, evaluating whether the model can adopt detector-agnostic GW features rather than memorising detector-specific noise.
@@
We first evaluate zero-shot inference by applying the model trained purely on ET noise directly to the LVK one, without any fine-tuning or weight updates. This results in a loss of predictive power, indicating that ET-trained features do not transfer to the LVK domain in a zero-shot setting. This is naturally explained by the model having learned a detector-conditioned latent space rather than an invariant one, and therefore, requiring domain adaptation through fine-tuning. Implementing transfer learning through fine tuning, the performance gain is significant compared to training from scratch across all inferred parameters, ranging between $9\%-47\%$ improvement for MAE (see Tables 4 and 5). The spin parameter shows the biggest improvements ($21-47\%$), suggesting that the pretrained model learned particularly robust representations of spin-related features that transfer effectively between detectors. The mass parameters and distance also show consistent improvements of $13-27\%$, indicating that the model successfully captures detector-agnostic physical relationships. The convergence speed results show that the pretrained model achieves the same performance targets as models trained from scratch in dramatically less time, up to $15\times$ faster. This speedup is particularly pronounced at lower SNR = 10, where the pretrained model reaches target performance in just 2-3 minutes compared to 30-50 minutes for scratch training. Interestingly, the speedup is somewhat reduced SNR = 30, possibly because cleaner signals allow scratch-trained models to converge more efficiently. These results collectively demonstrate that pretraining on one detector configuration creates transferable representations that not only improve final performance, but also significantly reduce the computational cost of adapting to new detector environments. We remind that while fine-tuning on LVK noise shows successful adaptation, both detectors still share broadly similar noise structure above 20 Hz. The more demanding task of adapting LVK-trained models to ET’s full 5 Hz range remains open for future work.
@@
To evaluate the generalisation of GraviBERT against waveform modelling, we performed a cross-waveform approximant validation. We first reconstructed the held-out test samples from the large dataset (originally generated with SEOBNRv4) using the IMRPhenomD approximant, adjusting the signals to match the temporal grid of the original SEOBNRv4 training data. Inference was then performed on these new waveforms using the frozen weights of our fully trained model (pretrained and trained on ET noise and SEOBNRv4 approximant) without any further fine-tuning, that is, at zero-shot inference. The results are shown in Table 6. While the baseline model achieves $R^{2}>0.93$ for all parameters at SNR = 10, the cross-waveform approximant inference appears to introduce a systematic bias. This effect is particularly evident for the effective spin parameter, where $R^{2}=0.65$ at SNR = 10. Spin is intimately connected with the subtle phase evolution during inspiral, which is modelled differently in SEOBNRv4 and IMRPhenomD. Nevertheless, the model retains predictive power for the dominant parameters $(m_{1},m_{2},d)$ at zero-shot inference, with $R^{2}$ remaining consistently between [0.80,0.97]. This suggests that the model has extracted representations of the signal envelope and frequency chirp, which are approximant-agnostic, while the specific phase alignment remains sensitive to the training domain.
@@ 5.4 Benchmarking
George & Huerta (2018) [27] demonstrated CNN-based parameter estimation achieving mean relative errors below 10% for mass parameters on real LVK data. Shen et al. (2022) [41] achieved 2 ms inference time using Bayesian neural networks trained on over 10 million waveforms for 4-parameter estimation, with relative errors of 7–12% for masses and 5–13% for remnant properties at SNR $\geq$ 15.
@@ 6 Conclusions and Outlook
We have introduced GraviBERT, a novel deep learning model for GW signal analysis that fuses an Inception-inspired multi-scale convolutional frontend with a transformer encoder. The key novelty of GraviBERT lies in its self-supervised pretraining strategy, which enables learning of transferable representations of the underlying physics, followed by fine-tuning on labeled data. This is facilitated by the feature extractor, which learns local temporal features, and the transformer’s self-attention mechanism, which captures global dependencies. In the current work, we demonstrate its capabilities using a regression head that outputs point estimates for source parameters backed by MC dropouts.
Our two-stage training strategy—leveraging BERT-style pretraining followed by targeted fine-tuning—yields substantial improvements in both training efficiency and inference accuracy, while demonstrating promising adaptation to new detector configurations and waveform approximants. When evaluated on in-domain data (ET), pretraining reduces the MAE by up to $31\%$ and accelerates training convergence by up to $6\times$. Furthermore, the mean relative precision of the inferred masses and distances reaches the few-percent level, while the MAE in effective spin is as low as $\sim 10^{-3}$ for the challenging case of SNR = 10555These results were obtained without exhaustive hyperparameter optimisation.. We note that the reported distance corresponds to an effective luminosity distance rescaled to fixed SNR, rather than the actual astrophysical luminosity distance.
The advantages are even more pronounced in domain adaptation from ET to LVK data. By fine-tuning on a small target dataset on LVK data, the pretrained model on ET noise converges up to $15\times$ faster and achieves an error reduction of up to $47\%$ compared to a model trained from scratch on LVK. This enhanced predictive precision is particularly notable in the low SNR regime, while the flexibility of the model’s architecture paves the way for broad applicability across multi-messenger astronomy. We note that while our estimation of inherent model uncertainty relied on the probabilistic approach of MC Dropouts, our parameter inference itself was not Bayesian. A Bayesian treatment could be pursued by replacing the regression head with normalising flows.
@@
The motivation for developing foundation-style modelling lies in the ability to generalise across a wide range of tasks and datasets, serving as adaptable tools for diverse downstream applications. Foundation models have transformed fields such as natural language processing by enabling rapid fine-tuning and transfer learning. For GW astrophysics, a similar approach could accelerate scientific discovery by enabling fast, accurate, and robust inference as new detectors, noise environments, and physical scenarios emerge. In this regard, GraviBERT represents a first step forward towards establishing a pretrained baseline framework for GW astronomy, which can be subsequently fine-tuned efficiently.
Natural extensions of the present work include applying GraviBERT to full Bayesian parameter estimation, inference of the complete set of source parameters for binary mergers, and comprehensive hyperparameter optimisation — all of which follow directly from the current framework. Additionally, our modelling could be expanded to cover neutron star mergers, alternative waveform approximants, and multi-channel detector configurations such as the full triangular ET design. More broadly, the architecture and pretraining approach presented here could in principle be extended to account for multi-messenger data, and to probing deviations from GR through their imprint on the waveform amplitude and phase evolution [9, 40, 6, 8]. On the architectural side, enhancements such as frequency-domain inputs, variable sampling rate, and comprehensive hyperparameter optimization could further improve performance and generality. We leave these directions for future work.
Acknowledgements & code
We thank François Charton, Nikos Karnesis, Fillipos Kokkinos, Nick Stergioulas and David Trestini for discussions and feedback. The authors were funded by the Czech Grant Agency (GAČR) under the grant number 21-16583M. The bulk of our computations were carried out at the MetaCentrum cluster in Czechia. The code and pretrained models are publicly available at this link.
@@
Table 2: Per-parameter metrics across different dataset sizes and SNR values for the model trained from scratch (not pretrained) and the pretrained and fine-tuned model, evaluated on the held-out test set with size of $10,000$ samples. The same detector configuration is kept throughout (ET). The reported metrics are measured on the mean values obtained from an ensemble of $100$ stochastic forward passes using MC Dropout for each test sample. We remind that $d$ corresponds to an effective luminosity distance rescaled to fixed SNR. For each metric, the best value across the two training configurations is shown in bold. The results are rounded to two decimal points.
@@
Table 3: Relative MAE improvement from pretraining followed by fine-tuning, compared to training from scratch, according to the results of Table 2. Positive values indicate the percentage reduction in error achieved through pretraining. The gains are calculated with original unrounded numbers.
@@
Table 4: Domain adaptation performance of the pretrained model evaluated with MAE. The comparison is between the model pretrained on ET noise and fine-tuned on LVK noise waveforms, and a model trained on LVK noise waveforms from scratch. The clean waveforms used cross both datasets were generated using the same procedure, with only the detector noise differing between domains. The results are rounded to two decimal points. The gains are calculated with original unrounded numbers.
@@
Table 5: Same as Table 4, but comparing the convergence speed between the model pretrained on ET noise and fine-tuned on LVK noise waveforms, and a model trained on LVK noise waveforms from scratch. As before, the “Target Loss" is the best validation loss achieved by the corresponding model trained from scratch and the speedup factor quantifies how many times faster the pretrained model reaches this same performance target.
@@
Table 6: Zero-shot performance of the model on a new waveform approximant at fixed detector noise (ET), quantified using MAE and $R^{2}$. The model was pretrained and subsequently trained on SEOBNRv4. It can be seen that zero-shot inference (without fine-tuning) already yields encouraging performance in the cross-waveform approximant setting (see also Section 5.3). The results are rounded to two decimal points.
@@
Table 7: Domain adaptation performance of the pretrained model on a new waveform approximant at fixed detector noise (ET), evaluated with MAE. The comparison is between the model pretrained on SEOBNRv4 waveforms and fine-tuned on IMRPhenomD waveforms, and a model trained on IMRPhenomD waveforms from scratch. The results are rounded to two decimal points. The gains are calculated with original unrounded numbers.
@@
7 Appendix
In this Appendix we present training loss curves and a table overviewing the hyperparameter choice for our analysis.
(a) Pretraining loss curve as a function of epochs.
(b) Val loss as a function of cumulative training time for models trained from scratch (orange) and fine-tuned (blue), across all configurations. The dotted horizontal lines (orange, blue) and the associated bullet points indicate the target loss. We refer to Table 1 for detailed numerical values.
Figure A1: Training dynamics during pretraining, and fine-tuning. During pretraining, training and validation losses decrease rapidly during the early stages of training, followed by a gradual increase at later epochs. This behaviour is consistent with the non-stationary nature of the self-targeting pretraining objective, where the reconstruction target is generated by the same evolving network.
Table A1: Hyperparameter configuration for the pretraining and training stages of the model.